OpenStudy (anonymous):

If you doubled the radius of a sphere, how would the volume change? A. The volume would double. B. The volume would triple. C. The volume would quadruple. D. The volume would increase to be eight times that of the original volume.

OpenStudy (anonymous):

The volume of a sphere is defined as: V=(4*pi*r^3)/3 then, if we double the radius of the sphere (r=2r): V=(4*pi*(2r)^3)/3 V=(4*pi*(2^3*r^3))/3 V=(4*pi*r^3)/3*8

OpenStudy (anonymous):

This would mean that by doubling the radius of the sphere the volume would increase eight times that of the original volume